How do I graph the problem f(x)=|x+2|-5 also how do I find the range, vertex, and x and y intercepts?
Answer by josgarithmetic(39838)

Understand at which x the function inside the absolute value becomes zero. At x=-2.

When x<-2, x+2<0
so y=-(x+2)-5
y=-x-7

When x=-2, y=-5

When x>-2, x+2>0
so y=x+2-5
y=x-3